# Belief Scene Graphs: Expanding Partial Scenes with Objects through Computation of Expectation

### v1

@@ Abstract
In this article, we propose the novel concept of Belief Scene Graphs, which are utility-driven extensions of partial 3D scene graphs, that enable efficient high-level task planning with partial information. We propose a graph-based learning methodology for the computation of belief (also referred to as expectation) on any given 3D scene graph, which is then used to strategically add new nodes (referred to as blind nodes) that are relevant for a robotic mission. We propose the method of Computation of Expectation based on Correlation Information (CECI), to reasonably approximate real Belief/Expectation, by learning histograms from available training data. A novel Graph Convolutional Neural Network (GCN) model is developed, to learn CECI from a repository of 3D scene graphs. As no database of 3D scene graphs exists for the training of the novel CECI model, we present a novel methodology for generating a 3D scene graph dataset based on semantically annotated real-life 3D spaces. The generated dataset is then utilized to train the proposed CECI model and for extensive validation of the proposed method. We establish the novel concept of Belief Scene Graphs (BSG), as a core component to integrate expectations into abstract representations. This new concept is an evolution of the classical 3D scene graph concept and aims to enable high-level reasoning for the task planning and optimization of a variety of robotics missions. The efficacy of the overall framework has been evaluated in an object search scenario, and has also been tested on a real-life experiment to emulate human common sense of unseen-objects.
@@ I INTRODUCTION
The growing need to develop high-level autonomy for robotic systems has set a broad range of challenges for different focus areas in the robotics research community. One of the most crucial challenge is the development of robotic capabilities that will enable the robots to plan their actions dynamically in the presence of uncertainty. Recently proposed planners incorporate the concept of belief for the task of navigation [1], variability estimation [2] and [3] localization of objects. However, this exciting research area is still in its initial stage and the majority of these robotic missions lack the capability to reason and extract higher level information from the scene, based on partial and uncertain information, as is the case with belief (also referred to as expectation in the rest of this article).
@@
In order to illustrate this utility of expectation in the context of robotic mission planning, let us consider the following example: if we were to leave a human and a robot (e.g. a service robot) in a completely unexplored area (e.g. a house with multiple stories) and ask them to make the laundry, the first task will be to look for dirty clothes followed by the washing machine and eventually the detergent. In this scenario, the expected behavior from the robot is to extensively look for such a set of objects, while exploring the totality of the building. On the other hand, a human being would take a fast look into each of the rooms to search for the one which is most likely to contain the objects (e.g. bedroom and washing room). Moreover, a human being may also look for the respective rooms in specific areas of the building, despite not having fully explored the building. The reasoning behind this decision is based on expectation and could be expressed as: I need to find dirty clothes, that are likely to be in the bedroom. The bedroom is probably on the second floor, so I must look for the stairs, which may be close to the entrance door.
@@
To enable reasoning and planning in such a complex mission, recently 3D scene graphs (3DSG) have been introduced and are gaining popularity [4]. In the context of 3D scene graphs, we can understand expectation as the likelihood of finding a specific object given a series of objects that have already been found. Substantial research has already been done in this field to improve and optimize the generation of 3D scene graphs, as well as for their implementation in a broader range of robotic tasks [5]. Nevertheless, even in the most recent state-of-the-art frameworks [6, 7, 8, 9, 10, 11, 12, 13, 14], the concept of 3D scene graphs remains pretty much unchanged from when it was first proposed [15]. In contrast to previous works, in this article we present an evolution of the classical concept, namely Belief Scene Graphs, with the aim of enabling high-level reasoning for task planning and optimization of different robotic missions, given a partial representation of the environment in the form of a 3D scene graph. An overview of our approach is presented in Fig. 1.
@@
Figure 1: Depiction of a high level task planning process based on Belief Scene Graphs, where the environment is first abstracted into data through perception in order to build a 3D scene graph representing building, rooms and objects, and then transform into the set of histograms $\textbf{E}^{\prime}_{\mathcal{N}}(C)\times O^{\prime}_{\mathcal{N}}$ that are input to the proposed CECI model to output the predicted set of histograms $\textbf{E}^{\prime\prime}_{\mathcal{N}}(C)\times O^{\prime\prime}_{\mathcal{N}}$ used to build a Belief Scene Graph with Blind Nodes and to draw expectation about the location of an Artifact, which enables a sequential high-level task planning.
I-A Related Work
@@
3D Scene graphs have emerged as a popular alternative for the encoding of scenes into intuitive and manageable data representations [6, 7, 8, 9, 10, 11, 12]. 3D scene graphs encode the environment in an abstract representation comprising of several layers (e.g. buildings, rooms and objects), in the form of nodes and edges, where nodes represent objects and edges represent the relationships among the nodes. Among the first works to propose the notion of 3D scene graphs is [15], where a semi-automatic framework for the construction of 3D scene graphs is presented. Since then, works such as in [13] and [14] have proposed frameworks for the optimized construction of 3D dynamic scene graphs in more complex large indoor environments. Nevertheless, it is only in the most recent state-of-the-art that we find the introduction of the idea of belief/expectation into 3D Scene Graphs, with the intention of enabling deeper reasoning with just partial information. Among these works we can find [3], where a spatial commonsense graph is proposed in order to solve the problem of localization of objects in partial 3D scans. Furthermore, [2] proposed a variable scene graph for semantic scene variability estimation, where three varieties of semantic scene change are identified: changes in position, semantic state, or scene composition. Finally, [1] presents a semantic belief graph for semantic-based planning under perceptual uncertainty to address navigation in extreme environments.
I-B Contributions
@@
In contrast to the state-of-the-art, the contributions of this article are summarized as follows. In this work, the focus is on extending the capability of a given partial 3D scene graph $\mathcal{G}^{\prime}$, in which objects of a set $C$ relevant for a robotic mission are missing, by proposing an approach to meaningfully incorporate $C$ as nodes into the scene graph $\mathcal{G}^{\prime}$. Towards this end, we propose a novel graph-based learning methodology, for the Computation of Expectation of finding objects in $C$ based on Correlation Information (CECI) between $C$ and the objects that are part of the given 3D scene graph $\mathcal{G}^{\prime}$. We design and implement a novel Graph Convolutional Network (GCN) model for CECI of 3D scene graphs. As no repository of 3D scene graphs is available for the training of the novel CECI model, we present a novel methodology for utilizing semantically annotated real-life 3D spaces to generate a 3D scene graphs dataset. The generated dataset is then utilized to train the proposed CECI model and also to validate the proposed methodology. Secondly, we establish the novel concept of Belief Scene Graphs (BSG), as a utility-enhanced extension of a given incomplete scene graph $\mathcal{G}^{\prime}$, by incorporating objects in $C$ into $\mathcal{G}^{\prime}$, using the learnt CECI information. Belief Scene Graphs enable high-level reasoning and optimized task planning involving set $C$, which was impossible with the incomplete $\mathcal{G}^{\prime}$. Finally, through extensive evaluation, we conclusively demonstrate the enhanced performance of the proposed BSG over classical 3DSG for the task of single-object and multi-object search and also present real-world results for prediction of unseen-objects by constructing a Belief Scene Graphs using information gathered by a legged robotic platform through the exploration of an unknown indoor environment.
@@ II PROBLEM FORMULATION
Modeling Expectation. Expectation can be intuitively modelled as the conditional probability of an object $a$ to be found given an object $b$ has already been found. In the context of 3D scene graphs, the probability for $a$ to be found depends on a set of $n\in\mathbb{Z^{+}}$ objects and room observations $\textbf{B}=\{b_{1},...,b_{n}\}$ which are assumed to be independent of each other. Then the expectation for an object $a$ to be found, given a series of objects and rooms B that had already been found, is given by: $\boldsymbol{P(a|B)}=\sum_{i=1}^{n}P(a\cap b_{i})/\prod_{i=1}^{n}P(b_{i})$. Note that the object $a$ and observation $b_{i}$ are being treated as singleton events with associated probabilities. In addition, in several robotic applications, we would like to know the probability of finding an object in a specific area or room, in other words, the expectation of finding $a$ in a specific node $v_{s}\in\mathcal{V}$ of the set of nodes $\mathcal{V}=\{v_{1},...,v_{w}\}$ with $w\in\mathbb{Z^{+}}$, or $\boldsymbol{P(a|B_{v_{s}})}$, where now we have the set of observations $\textbf{B}_{v_{s}}$ with respect to the node ${v_{s}}$. Furthermore, we would like to estimate the probability of finding not only a single object but also a set of objects $\textbf{A}=\{a_{1},...,a_{m}\}$, that is
$\boldsymbol{D_{v_{s}}(A)}=\begin{pmatrix}P(a_{1}|B_{{v_{s}}})&P(a_{2}|B_{{v_{s}}})&\cdots&P(a_{m}|B_{{v_{s}}})\end{pmatrix}$ (1)
where $\textbf{D}_{v_{s}}$ is the set of probabilities of the expected objects in the set A over the node ${v_{s}}$.
@@
Many robotic applications also require information about the attributes of the expected objects (e.g. size and number of objects). Let us assume now the set of attributes $\textbf{F}_{b_{i}}=\{f_{b_{i}1},...,f_{b_{i}u}\}$ for each observation $b_{i}\in\textbf{B}$. In this scenario, we need to estimate the conditional probability of the object $a_{j}$ with respect to the observation $b_{i}$ with the set of attributes $\textbf{F}_{b_{i}}$ as:
$\boldsymbol{P(a_{j}|\{b_{i},F_{b_{i}}\})}=\frac{P(a_{j}\cap b_{i}\cap f_{b_{i}1}...\cap f_{b_{i}u})}{P(b_{i}\cap f_{b_{i}1}...\cap f_{b_{i}u})}$ (2)
@@
Likewise, the conditional probability of an object $a_{j}$ with a specific attribute $f_{a_{j}k}\in\textbf{F}_{a_{j}}$ can be estimated through:
$\boldsymbol{P(f_{a_{j}k}|\{b_{i},F_{b_{i}}\})}=\frac{P(f_{a_{j}k}\cap b_{i}\cap f_{b_{i}1}...\cap f_{b_{i}u})}{P(b_{i}\cap f_{b_{i}1}...\cap f_{b_{i}u})}$ (3)
where $f_{a_{j}k}$ is the $k^{\text{th}}$ element of the set $\textbf{F}_{a_{j}}$.
@@
This reasoning is similar to the way human beings draw expectations. For a human, the expectation to find an object $a_{j}$ in a given room $v_{s}$ is based on the nature of the object and the room itself (i.e. $\textbf{F}_{a_{j}}$ and $\textbf{F}_{{v_{s}}}$), the objects found in previous rooms (i.e. $\textbf{B}_{v_{s}}$) and their similarity with the expected object $a_{j}$. Even the nature of the building itself may temper the expectation of finding a specific object, for example, we will have lower expectations to find a pillow in an office building than in a residential building.
@@
CECI. Let $\textbf{D}_{\mathcal{V}}$ denote the real set of conditional probabilities of an object set A with respect to all the nodes $\mathcal{V}$ of a given 3D scene graph. In real-life, translating the intuitive idea of expectations into a computational method for deriving $\textbf{D}_{\mathcal{V}}$ is challenging, due to the heavy data requirements for the purpose. In order to overcome this challenge, in this article, we look to find a reasonably approximate set of probability distributions $\textbf{E}_{\mathcal{V}}$ for the set of probabilities $\textbf{D}_{\mathcal{V}}$, from the histograms of the set of nodes $\mathcal{V}$, where the histogram is equal to $\textbf{E}_{\mathcal{V}}\times O_{\mathcal{V}}$, where $O_{\mathcal{V}}\in\mathbb{Z}^{+}$ is the number of object nodes connected to each respective node in $\mathcal{V}$. Due to the sheer variability amongst real-life applications, the histogram of any given room is either unknown or hard to estimate for a field robot online. This is due to multiple factors, such as:
@@
  • Time. The exploration is time-intensive or the robotic platform is time constrained (e.g. drones flight time).
@@
  • Blockades. The physical limitations of the platform do not allow for the exploration of specific areas (e.g. a mobile robot reaching an steep change on elevation).
@@
  • Observability. The platform can only partially perceive the scene (e.g. a drone looking through a window).
@@
In order to be able to estimate $\textbf{E}_{\mathcal{V}}$ in partially known environments, and to allow the scalability of the framework, we propose a new model of Graph Convolutional Netowrks (GCN) for Computation of Expectation base on Correlation Information (CECI) of 3D scene graphs. The novel CECI model aims to learn the underlying probability distribution for each node based on a dataset of sample environments. The CECI model should then be able to use sparse information about any given 3D scene graph topology, to draw expectations about possible objects present in the environment. The next step is then to define the problem in terms of a GCN architecture, where the input is any given sparse 3D scene graph $\mathcal{G}^{\prime}$ with a set of histograms $\textbf{E}^{\prime}_{\mathcal{N}}(C)\times O^{\prime}_{\mathcal{N}}$ as node attributes over a set of labeled object classes $C$, where $\mathcal{N}=\{n_{1},\dots,n_{Q}\}$ is the set of nodes of interest, and the output is the set of histograms $\textbf{E}^{\prime\prime}_{\mathcal{N}}(C)\times O^{\prime\prime}_{\mathcal{N}}$ which can be used to approximate $\mathcal{G}^{\prime}$ to the ground truth graph $\mathcal{G}$ with a set of histograms $\textbf{E}_{\mathcal{N}}(C)\times O_{\mathcal{N}}$, as depicted in Fig. 1.

### v2

@@ Abstract
In this article, we propose the novel concept of Belief Scene Graphs, which are utility-driven extensions of partial 3D scene graphs, that enable efficient high-level task planning with partial information. We propose a graph-based learning methodology for the computation of belief (also referred to as expectation) on any given 3D scene graph, which is then used to strategically add new nodes (referred to as blind nodes) that are relevant to a robotic mission. We propose the method of Computation of Expectation based on Correlation Information (CECI), to reasonably approximate real Belief/Expectation, by learning histograms from available training data. A novel Graph Convolutional Neural Network (GCN) model is developed, to learn CECI from a repository of 3D scene graphs. As no database of 3D scene graphs exists for the training of the novel CECI model, we present a novel methodology for generating a 3D scene graph dataset based on semantically annotated real-life 3D spaces. The generated dataset is then utilized to train the proposed CECI model and for extensive validation of the proposed method. We establish the novel concept of Belief Scene Graphs (BSG), as a core component to integrate expectations into abstract representations. This new concept is an evolution of the classical 3D scene graph concept and aims to enable high-level reasoning for task planning and optimization of a variety of robotics missions. The efficacy of the overall framework has been evaluated in an object search scenario, and has also been tested in a real-life experiment to emulate human common sense of unseen-objects.
For a video of the article, showcasing the experimental demonstration, please refer to the following link: https://youtu.be/hsGlSCa12iY
@@ I INTRODUCTION
The growing need to develop high-level autonomy for robotic systems has set a broad range of challenges for different focus areas in the robotics research community. One of the most crucial challenge is the development of robotic capabilities that will enable the robots to plan their actions dynamically in the presence of uncertainty. Recently proposed planners incorporate the concept of belief for the task of navigation [1], variability estimation [2] and [3] localization of objects. However, this exciting research area is still in its initial stage and a majority of such robotic missions lack the capability to reason and extract higher-level information from the scene, based on partial and uncertain information, as is the case with belief, also referred to as expectation in the rest of this article.
@@
In order to illustrate the utility of expectation in the context of robotic mission planning, let us consider the following example: if we were to leave a human and a robot in a completely unexplored area and ask them to find a specific object, the expected behavior from the robot is to extensively look for such an object, while exploring the totality of the building. On the other hand, a human would take a fast look into each of the rooms to search for the one that is most likely to contain the object. Moreover, a human may also look for the respective rooms in specific areas of the building, despite not having fully explored the building. The reasoning behind this decision is based on expectation and could be expressed as: I need to find a pillow, that is likely to be in the bedroom, the bedroom is probably on the second floor, so I must look for the stairs, which may be close to the entrance door.
@@
To enable reasoning and planning in such a complex mission, 3D scene graphs (3DSG) have been introduced recently and are gaining popularity [4]. In the context of 3D scene graphs, we can understand expectation as the likelihood of finding a specific object given a series of objects that have already been found. Substantial research has already been done in this field to improve and optimize the generation of 3D scene graphs, as well as for their implementation in a broad range of robotic tasks [5]. Nevertheless, even in the most recent state-of-the-art frameworks [6, 7, 8, 9, 10, 11, 12, 13, 14], the concept of 3D scene graphs remains pretty much unchanged from when it was first proposed [15]. In contrast to previous works, in this article, we present an evolution of the classical concept, namely Belief Scene Graphs, with the aim of enabling high-level reasoning for task planning and optimization of different robotic missions, given a partial representation of the environment in the form of a 3D scene graph. An overview of our approach is presented in Fig. 1.
@@
Figure 1: Depiction of a high-level task planning process based on Belief Scene Graphs, where the environment is first abstracted into data through perception, in order to build a 3D scene graph representing building, rooms and objects, and then transform into the set of histograms $E^{\prime}_{\mathcal{N}}(C)\times O^{\prime}_{\mathcal{N}}$ that are input to the proposed CECI model to output the predicted set of histograms $E^{\prime\prime}_{\mathcal{N}}(C)\times O^{\prime\prime}_{\mathcal{N}}$ used to build a Belief Scene Graph with blind nodes and to draw expectation about the location of an artifact, which enables a sequential high-level task planning.
— Related Work
@@
3D Scene graphs have emerged as a popular alternative for the encoding of scenes into intuitive and manageable data representations [6, 7, 8, 9, 10, 11, 12]. 3D scene graphs encode the environment in an abstract representation comprising of several layers (e.g. buildings, rooms and objects), in the form of nodes and edges, where nodes represent objects and edges represent the relationships among the nodes. Among the first works to propose the notion of 3D scene graphs is [15], where a semi-automatic framework for the construction of 3D scene graphs is presented. Since then, works such as  [13] and [14] have proposed frameworks for the optimized construction of 3D dynamic scene graphs in more complex and large indoor environments.
Nevertheless, it is only in the most recent state-of-the-art that we find the introduction of the idea of belief/expectation into 3D Scene Graphs, with the intention of enabling deeper reasoning with just partial information. Among these works we can find [3], where a spatial commonsense graph is proposed in order to solve the problem of localization of objects in partial 3D scans. Furthermore, [2] proposed a variable scene graph for semantic scene variability estimation, where three varieties of semantic scene change are identified: changes in position, semantic state, or scene composition. Finally, [1] presented a semantic belief graph for semantic-based planning under perceptual uncertainty to address navigation in extreme environments.
— Contributions
@@
In contrast to the state-of-the-art, the contributions of this article are summarized as follows. In this work, the focus is on extending the capability of a given partial 3D scene graph $\mathcal{G}^{\prime}$, in which objects of a set $C$ relevant for a robotic mission are missing, by proposing an approach to meaningfully incorporate $C$ as nodes into the scene graph $\mathcal{G}^{\prime}$. Towards this end, we propose a novel graph-based learning methodology, for the Computation of Expectation of finding objects in $C$ based on Correlation Information (CECI) between $C$ and the objects that are part of the given 3D scene graph $\mathcal{G}^{\prime}$. We design and implement a novel Graph Convolutional Network (GCN) model for CECI of 3D scene graphs. As no repository of 3D scene graphs is available for the training of the novel CECI model, we present a novel methodology for utilizing semantically annotated real-life 3D spaces to generate a 3D scene graphs dataset. The generated dataset is then utilized to train the proposed CECI model and also to validate the proposed methodology. Secondly, we establish the novel concept of Belief Scene Graphs (BSG), as a utility-enhanced extension of a given incomplete scene graph $\mathcal{G}^{\prime}$, by incorporating objects in $C$ into $\mathcal{G}^{\prime}$, using the learnt CECI information. Belief Scene Graphs enable high-level reasoning and optimized task planning involving set $C$, which was impossible with the incomplete $\mathcal{G}^{\prime}$. Finally, through extensive evaluation, we conclusively demonstrate the enhanced performance of the proposed BSG over classical 3DSG for the task of single-object and multi-object search and also present real-world results for prediction of unseen-objects by constructing a Belief Scene Graph using information gathered by a legged robotic platform through the exploration of an unknown indoor environment.
@@ II PROBLEM FORMULATION
Modeling Expectation. Expectation can be intuitively modeled as the conditional probability of an object to be found given other objects had already been found. Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ denote a 3D Scene Graph, defined as a tuple of nodes $\mathcal{V}$ and edges $\mathcal{E}$. Similarly, let $C=\{c_{1},...,c_{n}\}$ with $n\in\mathbb{Z^{+}}$ denote the set of labeled object classes present in the graph $\mathcal{G}$. In this context, the probability for a given object class $c_{i}\in C$ to be found depends on a set of objects and room observations $B=\{b_{1},...,b_{m}\}$, $m\in\mathbb{Z^{+}}$, which are assumed to be independent of each other. Then the expectation for an object $c_{i}$ is given by $\boldsymbol{P(c_{i}|B)}=\sum_{j=1}^{m}P(c_{i}\cap b_{j})/\prod_{j=1}^{m}P(b_{j})$. Note that the object $a_{i}$ and observation $b_{j}$ are being treated as single events with associated probabilities. Likewise, we can estimate the set of probabilities for finding every object class in the set $C$, that is:
$\boldsymbol{D=}\begin{pmatrix}P(c_{1}|B)&P(c_{2}|B)&\cdots&P(c_{n}|B)\end{pmatrix}.$ (1)
@@
Furthermore, many robotic applications also require information about the number of expected objects, which is encoded into the attribute set $A_{B}=\{a_{b1},...,a_{bm}\}$, with each element corresponding to an observation in set $B$. In this scenario, we need to estimate the conditional probability of the object $c_{i}$ with respect to the observations $B$ with $A_{B}$ attributes as:
$\boldsymbol{P(c_{i}|B,A_{B})}=\sum_{j=1}^{m}P(c_{i}\cap b_{j}\cap a_{bj})/\prod_{j=1}^{m}P(b_{j}\cap a_{bj})$ (2)
@@
Likewise, the conditional probability of an object class $c_{i}$ with $a_{ci}\in A_{C}$ attribute can be estimated through:
$\boldsymbol{P(a_{ci}|B,A_{B})}=\sum_{j=1}^{m}P(a_{ci}\cap b_{j}\cap a_{bj})/\prod_{j=1}^{m}P(b_{j}\cap a_{bj})$ (3)
@@
This reasoning is similar to the way humans draw expectations. For a human, the expectation to find an object $c_{i}$ in a given room is based on the nature of the object and the room itself, the objects found in previous rooms (i.e. $B$) and their similarity with the expected object $c_{i}$. Even the nature of the building itself may temper the expectation of finding a specific object, for example, we will have lower expectations to find a pillow in an office building than in a residential building.
@@
Approximating Expectation. Let $D_{\mathcal{V}}$ denote the real set of conditional probabilities of a set of labeled object classes $C$ with respect to all the nodes $\mathcal{V}$ of a given 3D scene graph. In real-life, translating the intuitive idea of expectations into a computational method for deriving $D_{\mathcal{V}}$ is challenging, due to the heavy data requirements for the purpose. To overcome this challenge, in this article, we look to find a reasonably approximate set of probabilities $E_{\mathcal{V}}$ for the set of probabilities $D_{\mathcal{V}}$, from the histograms of the set of nodes $\mathcal{V}$, where the histogram is equal to $E_{\mathcal{V}}\times O_{\mathcal{V}}$, where $O_{\mathcal{V}}\in\mathbb{Z}^{+}$ is the number of object nodes connected to each respective node in $\mathcal{V}$. Due to the sheer variability amongst real-life applications, the histogram of any given room is either unknown or hard to estimate for a field robot online. This is due to multiple factors, such as:
@@
  • Time. The exploration is time-intensive or the robotic platform is time-constrained (e.g. drone flight time).
@@
  • Blockades. The physical limitations of the platform do not allow for the exploration of specific areas (e.g. a mobile robot reaching a steep change in elevation).
@@
  • Occlusion. The platform can only partially perceive the scene (e.g. a drone looking through a window).
@@
Computation of Expectation based on Correlation Information. In order to be able to estimate $E_{\mathcal{V}}$ in partially known environments, and to allow the scalability of the framework, we propose a new model of Graph Convolutional Networks (GCN) for the Computation of Expectation base on Correlation Information (CECI) of 3D scene graphs. The novel CECI model aims to learn the underlying probability distribution for each node based on a dataset of sample environments. The CECI model should then be able to use sparse information about any given 3D scene graph topology, to draw expectations about possible objects present in the environment. The next step is then to define the problem in terms of a GCN architecture, where the input is any given sparse 3D scene graph $\mathcal{G}^{\prime}$ with a set of histograms $E^{\prime}_{\mathcal{N}}(C)\times O^{\prime}_{\mathcal{N}}$ as node attributes over a set of labeled object classes $C$, where $\mathcal{N}$ is the set of nodes of interest, and the output is the set of histograms $E^{\prime\prime}_{\mathcal{N}}(C)\times O^{\prime\prime}_{\mathcal{N}}$ which can be used to approximate $\mathcal{G}^{\prime}$ to the ground truth graph $\mathcal{G}$ with a set of histograms $E_{\mathcal{N}}(C)\times O_{\mathcal{N}}$, as depicted in Fig. 1.